OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 2(x − 1) −12? x >or= −5 x

OpenStudy (precal):

you are missing the inequality before the -12 perhaps

OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 2(x − 1) >or= −12? x >or= −5 x <or =−5 −5 <or = x A number line with a closed circle on negative 5 and shading to the right

OpenStudy (precal):

 can you finish the problem?